ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 134, 2003

Adin, R. M. See A. Berele

Berele, A. and R. M. Adin Denominators for the Poincaré series of invariants of matrices with involution

93

Berman, S., Y. Gao and S. Tan A unified view of some vertex operator constructions

29

Flicker, Y. Z. and D. Zinoviev On the symmetric square. Unstable twisted characters

307

Gao, Y. See S. Berman

Glasner, E. and B. Weiss Quasifactors of ergodic systems with positive entropy

363

Kissin, E., V. I. Lomonosov and V. S. Shulman Implementation of derivations and invariant subspaces

1

Levin, M. B. and E. Merzbach Central limit theorems for the ergodic adding machine

61

Lomonosov, V. I. See E. Kissin

Lukeš, J., J. Malý, I. Netuka, M. Smrčka and J. Spurný On approximation of affine Baire-one functions

255

Malý, J. See J. Lukeš

Massart, D. On Aubry sets and Mather’s action functional  157

Merzbach, E. See M. B. Levin

Müller, P. F. X. Family of complemented subspaces in VMO and its isomorphic classification

289

Netuka, I. See J. Lukeš

Pila, J. Entire functions having a concordant value sequence  317

Ruelle, D. Perturbation theory for Lyapunov exponents of a toral map: Extension of a result of Shub and Wilkinson

345

Sela, Z. Diophantine geometry over groups II: Completions, closures and formal solutions

173

Shelah, S. Successor of singulars: Combinatorics and not collapsing cardinals ≤ κ in (< κ)-support iterations

127

Shulman, V. S. See E. Kissin

Smrčka, M. See J. Lukeš Spurný, J. See J. Lukeš Tan, S. See S. Berman Weiss, B. See E. Glasner Zinoviev, D. See Y. Z. Flicker